# SMILES & SELFIES have to go: Representation of Molecules via Algebraic Data Types

### v2

@@ Abstract
The Algebraic Data Type (ADT) can be used as a computational framework for molecular representation for the purpose of advancing tasks in cheminformatics. This can include generative modles in the context of Bayesian machine learning via probabilistic programming. The ADT that we put forward, implements the ‘Dietz’ representation for molecular constitution via multigraphs of electron valence information, and uses 3D coordinate data to provide stereochemical information, easily enabling the representation of complex molecular phenomena such as organometallics, multi-center bonds, delocalized electrons, and resonant structures. Unlike traditional string-based representations such as SMILES and SELFIES, the ADT is much more flexible, yet retains desirable qualities from type-safety to seamless integration with Bayesian Probabilistic Programming. An extensive criticism of both SMILES and SELFIES, in this article, is given, along with criticisms of the so-called Future of SELFIES. An open-source library implemented in Haskell demonstrates the ADT along with experimental extensions demonstrating its use in reaction modelling, group theoretic applications, and integration with LazyPPL, a lazy probabilistic programming library. Also provided as an extension is the ability to represent electronic structures, including shells, subshells, and orbitals, significantly expanding its representational scope compared to other representations in the literature. These features position the proposed ADT as a robust alternative to existing molecular representations, addressing limitations such as inadequate support for 3D information and syntactic invalidity while offering a platform for innovative cheminformatics research. Accompanying discussions about the meaning of a ‘representation’ are included. The fully functioning GitHub library can be found at https://github.com/oliverjgoldstein/chemalgprog.
@@ 1 Introduction
The choice of molecular representation directly affects cheminformatics tasks such as novel molecule discovery and molecular property prediction [1]. More expressive representation enables the exploration of larger regions of chemical space, both as input to and output from machine learning models. However, representational ability alone does not determine the effectiveness of a molecular representation in ML applications. Computational efficiency of optimisation algorithms, model choice and the amount of data provided are also important factors, along with the ease—or even possibility of integrating representations with techniques such as geometric deep learning [2] or Bayesian machine learning [3]. In cheminformatics, we believe that when it comes to molecular representation, “the medium is the message.”.  [^1] From Marshall McLuhan. More specifically, the context of communication decides the operations which the transported data supports.
@@
Despite decades of research, string-based representations such as SMILES [4] and SELFIES [5] remain the most common representations for ML tasks [6, 7, 8], despite well documented short-comings in representational ability and with syntaxes themselves. For instance, standard SMILES strings cannot represent complex bonding systems, resonance structures, or delocalized bonding, and do not include any 3D information, with only a limited ability to describe stereochemical features. In addition, not all SMILES strings correspond to valid molecules, meaning that the output of ML models is, in many cases, gibberish [5], and small changes in SMILES strings can result in large differences in molecular structure, making property prediction difficult. Attempts to build upon strings have led to a profusion of new grammars and syntaxes to represent molecular phenomena, a problem in itself, such as BIGSMILES or CURLYSMILES etc.
@@
While proposing future work to address the limitations of string-based representations such as SELFIES, [5] suggest developing a “molecular programming language”. Reasoning that strings can encode Turing-complete programming languages (e.g., Python source code), they suggest SELFIES could itself become a programming language, extended with a syntax that allows for recursion (for/while loops) and logical statements. Although this reasoning appears to conflate strings as a data structure with strings as a sequence of ASCII characters, we ask the question: if molecules can be represented as data within well established programming languages, we ask, why not simply use one?
In this paper, we critique contemporary computational representations, their application and challenges with ML, and some of the proposed solutions relating to string-based representations — we then present a solution with a novel framework: the Algebraic Data Type (ADT), written in the functional and typed language, Haskell.
@@
The ADT has features of the Dietz representation for constitution [9], using multigraphs of atomic valence information. It contains optional features such as 3D atomic coordinate data to provide functions such as conformation energy prediction and configuration of molecules, as well as quantum information in the form of orbitals, which is a first in the literature.
The ADT is able to represent complex bonding systems (e.g. organometallics), resonance structures, and delocalised bonding. We further demonstrate extending the ADT in modelling reaction networks, and by using the ADT as an instantiation of a mathematical group. We also construct a probabilistic model in LazyPPL that generates and scores molecules based on arbitrary metrics to show the ADT’s ability to be used with probabilistic programming.
@@
2 Forms of Molecular Representation
In this paper, we distinguish three different usages of the term “molecular representation”:
@@ 2 Forms of Molecular Representation
    Storage representations refer to the format for storage of molecular data on-disk, that is, without interpretation by a computer program.
@@
    Transmission representations relate to the way molecules are sent between cheminformaticians, and may depend on the medium of communication used.
@@
    Computational representations refer to how molecules are represented inside programming language or program, specifically, the data type or types which are used to represent the program.
@@
3 Contemporary Molecular Representations
In this paper we consider Molecular Graphs, Molecular Hypergraphs, the Dietz representaion (via multigraphs), string based representations and their numerous extensions, fingerprints, Variational Auto Encoders (VAEs), and computation tables.
@@ 3.3 Multigraphs
Multigraphs are not widely used as a standard computational representation in cheminformatics or chemical graph theory, despite the protestations of Dietz [9]. Existing tools and frameworks generally rely on labeled graphs or weighted graphs to encode bond multiplicities and types.
@@
A multiset (or bag) is a generalization of a set where elements are allowed to appear more than once. Formally, a multiset $M$ over a universe $U$ is defined as a function:
@@ 3.4 Dietz representation
The constitution of a structure is represented in set theoretic notation, with two sets. Our ADT includes type information, which prevents invalid molecules as will be discussed. The first set contains all atoms in the structure as triples, containing the number of unshared valence electrons of the atom, a numerical index, and the atomic symbol. The second set describes “bonding systems”, or bonding relationships between atoms, and contains pairs consisting of the number of shared valance electrons, and a set of edges representing bonds over which the electrons are shared.
@@
Delocalised bonding is represented by a bonding relationship containing multiple set of edges. For instance, the delocalised pi electrons of a benzene ring would be described by a bonding relationship of the form $(6,\{\{c_{1},c_{2}\},\{c_{2},c_{3}\},\{c_{3},c_{4}\},\{c_{4},c_{5}\},\{c_{5},c_{6}\},\{c_{6},c_{1}\}\})$, where $c_{i}$ is the numerical index of each carbon atom in the ring.
@@
There are two hydrogen atoms, labelled 1 and 2, each with zero unshared valance electrons in the atomic core, and a single covalent bond with two shared valance electrons.
@@
A major advantage of the Dietz representation is its general model for the representation of bonding. As a result, it is straightforward to describe complex bonding systems such as organometallics, with ferrocene and diborane being given as examples in the original paper. The lack of labeling requirements for delocalised bonding stands in contrast to string based representations such as SMILES, which has explicit rules in the grammar (a form of labeling) for aromatic carbon, and two rules for different cases of aromatic nitrogen.
@@
As Krenn [1] puts it:
The same applies to other weak interactions such as hydrogen bonds, raising important questions as to which interactions should actually be captured in a digital molecular representation as a “bond” (and which should not) and how to automatically detect them from a set of atomic coordinates, ultimately leading to a rather arbitrary distinction between bonded and non-bonded. To quote Democritus: “Nothing exists except atoms and empty space; everything else is only opinion”.
@@ 3.5 SMILES
As molecular graphs, SMILES inherit the limitations of this representational framework. While SMILES can represent aromaticity through specific syntax[4], it cannot encode delocalized bonding beyond such cases. Resonance and tautomerism must either be simplified into a single structure or require multiple distinct structures to represent each form.
@@
Attempts to canonicalize SMILES, and the proprietary nature of the Daylight canonicalisation algorithm, has led to inconsistencies between implementations and research teams [11], though the OpenSMILES initiative has mitigated this issue by providing a standardized grammar.
@@ 3.6 SELFIES
Table 1: Derivation rules of SELFIES for molecules in the QM9 dataset. Table formatted to fit within the page.
@@
Furthermore, E-Z stereochemistry cannot provide accurate dihedral or torsional angles [27]. Proteins rely on these angles for many tasks, which may be critical for machine learning tasks [28]. As Dayalan explains ([21]) “Dihedral angles are of considerable importance in protein structure prediction as they define the backbone of a protein, which together with side chains define the entire protein conformation.”. Angles influence the energy associated with a conformation of a drug which in turn relates to the utility and stability of the drug [29]. Although SELFIES solves the syntactic invalidity problem of SMILES, many challenges remain.
3.6.1 The future of SELFIES and of molecular representation - Extensions of SELFIES and future projects
@@ 3.8 Barriers to a general, digital, molecular representation
The lack of quantum chemical information, which is not currently part of any modern digital representation, is also a barrier to representation, as expresed in the Kenn paper:
@@ 3.10 Impact on machine learning models and empirical verification
3.11 The Algebraic Data Type (ADT) as an Alternative
We believe that the ADT later introduced will solve all of the above mentioned problems, for representation, and for the impact on machine learning models. and will represent the general, digital molecular representation searched for, by moving beyond string-based paradigms of SMILES, SELFIES, and their numerous extensions.
@@ 3.12 Extensions of SMILES
New syntaxes or grammars for string-based representation have been introduced in an attempt to represent different kinds of molecules not representable with only SMILES. For instance, there are many syntaxes from just one vendor called Daylight222www.daylight.com.
@@
[33] introduces Balsa, a fully-specified language designed as a subset of SMILES. It aims to address the lack of a comprehensive SMILES specification and the resulting practical problems, such as the diversity of implementations, difficulties in standardization efforts, and challenges in developing extensions. Balsa provides a formal, machine-readable grammar that defines its syntax and describes semantics at three levels: constitution, configuration, and conformation. However, while Balsa improves specificity and compatibility, its compact representation may sacrifice some readability compared to more verbose alternatives. Additionally, like SMILES, Balsa strings can become complex and difficult to interpret for larger, more intricate molecules.
3.13 Fingerprints

### v3

# Representation of Molecules via Algebraic Data Types : Advancing Beyond SMILES & SELFIES

@@ Abstract
We introduce a novel molecular representation through Algebraic Data Types (ADTs) - composite data structures formed through the combination of simpler types that obey algebraic laws. By explicitly considering how the datatype of a representation constrains the operations which may be performed, we ensure meaningful inference can be performed over generative models (programs with sample and score operations). This stands in contrast to string-based representations where string-type operations may only indirectly correspond to chemical and physical molecular properties, and at worst produce nonsensical output.
The ADT presented implements the Dietz representation for molecular constitution via multigraphs and bonding systems, and uses atomic coordinate data to represent 3D information and stereochemical features. This creates a general digital molecular representation which surpasses the limitations of the string-based representations and the 2D-graph based models on which they are based. In addition, we present novel support for quantum information through representation of shells, subshells, and orbitals, greatly expanding the representational scope beyond current approaches, for instance in Molecular Orbital theory.
The framework’s capabilities are demonstrated through key applications: Bayesian probabilistic programming is demonstrated through integration with LazyPPL, a lazy probabilistic programming library; molecules are made instances of a group under rotation, necessary for geometric learning techniques which exploit the invariance of molecular properties under different representations; and the framework’s flexibility is demonstrated through an extension to model chemical reactions.
After critiquing previous representations, we provide an open-source solution in Haskell – a type-safe, purely functional programming language.
@@ 1 Introduction
The choice of molecular representation directly affects cheminformatics tasks such as novel molecule discovery and molecular property prediction [1]. More expressive representation enables the exploration of larger regions of chemical space, both as input to and output from machine learning models. However, representational ability alone does not determine the effectiveness of a molecular representation in ML applications. Computational efficiency, model convergence, and predictive accuracy are also important factors, along with the ease—or even possibility of integrating representations with techniques such as geometric deep learning [2] or Bayesian machine learning [3]. In cheminformatics, we believe that when it comes to molecular representation, “the medium is the message.”
@@
Despite decades of research, string-based representations such as SMILES [4] and SELFIES [5] remain the most common representations for ML tasks [6, 7], despite well documented short-comings in representational ability and with syntaxes themselves. For instance, standard SMILES strings cannot represent complex bonding systems, resonance structures, or delocalized bonding, and do not include any 3D information, with only a limited ability to describe stereochemical features. In addition, not all SMILES strings correspond to valid molecules, meaning that the output of ML models is, in many cases, gibberish [5], and small changes in SMILES strings can result in large differences in molecular structure, making property prediction difficult. Attempts to build upon strings have led to a profusion of new grammars and syntaxes to represent molecular and chemical phenomena, a problem in itself.
@@
While proposing future work to address the limitations of string-based representations such as SELFIES, Krenn et al (2022) [5] suggest developing a “molecular programming language”. Reasoning that strings can encode Turing-complete programming languages (e.g., Python source code), they suggest SELFIES could itself become a programming language, extended with a syntax that allows for recursion (for/while loops) and logical statements. Although this reasoning appears to conflate strings as a data structure with strings as a sequence of ASCII characters, we ask the question: if molecules can be represented as data within programming languages, we ask, why not simply use one?
In this article, we critique the string-based representations SMILES and SELFIES, along with their extensions, evaluating their structure, limitations, and suggested improvements. We assess molecular fingerprints, focusing on their role in similarity searching and cheminformatics workflows, and examine their applicability and challenges in machine learning contexts. Additionally, we evaluate how molecules are represented within programming environments, specifically considering RDKit as a representation, prior attempts to formalize molecular representations using Inductive Logic Programming (ILP), and previous work integrating functional programming and cheminformatics. We explore the mathematical primitives underlying these representations—vectors, graphs, and strings, including as a data type, and their implications for computational workflows. We consider existing representations in terms of their applicability to Bayesian Machine learning and Geometric Deep Learning, as well as reaction modeling. Finally, we conclude by presenting the ADT framework, implemented in the functional and strongly typed language, Haskell.
@@
The ADT that we introduce uses the Dietz representation for constitution of molecules[8], using multigraphs of atomic valence information. It contains optional features such as 3D atomic coordinate data to provide functions such as conformation energy prediction and configuration of molecules, as well as quantum information in the form of orbitals.
The ADT is a general model for molecular representation, surpassing the limitations of 2D-graph based representations. It can represent complex bonding systems, such as the delocalised and multi-centre bonding within organometallics, and through atomic coordinates it can represent geometry beyond a limited range of labeled stereochemical features.
We demonstrate the flexibility of the ADT by showing an extension to model chemical reactions, and beyond representation, we demonstrate the ADTs suitability for Bayesian Machine Learning techniques with a probabilistic model written with Lazy PPL, a probabilistic programming library. We also demonstrate that ADTs can be made groups under permutation of atomic index, a necessary requirement for integration with Geometric Deep Leaning Techniques.
@@
The length of this article is reflective of the cross-disciplinary nature of the work, and of the requirement to explain concepts from functional programming and type-theory that may not be necessary for a more familiar audience. Although other work has linked functional programming and cheminformatics before [9, 10], this work was primarily focused on implementing existing cheminformatics workflows within functional programming languages, and the benefits of immutable state and type-saftey. Instead, we provide a molecular representation, and the surrounding context for the importance of considering data types when chosing a representaion.
@@
1.1 Forms of Molecular Representation
In this article, we distinguish three different usages of the term “molecular representation”:
@@ 1.1 Forms of Molecular Representation
    Storage representations refer to the format for storage of molecular data.
@@
    Transmission representations refer to the way molecules are sent between cheminformaticians, and depend on the medium of communication used.
@@
    Computational representations refer to how molecules are represented inside of a programming language or program, specifically, the data type or types which are used to represent the molecule.
@@
1.2 Data Types
Before continuing with the article, we need to define some concepts used throughout:
Definition 1 (Data Type).
A data type classifies and structures the values a program can manipulate, describing how they are constructed, stored, and what operations are valid on them. In many programming languages, a data type encapsulates both the possible shapes (or “constructors”) of the data and the rules for operating on those shapes.
In Haskell, a data type is introduced with the data keyword and defines one or more constructors that specify exactly how to build valid values of that type. This not only ensures consistency (i.e., that every value has a well-defined form) but also guides how functions can safely pattern-match and handle each variant of the type.
Definition 2 (Typeclass).
A typeclass (in languages such as Haskell) is a mechanism for defining a set of operations or methods that one or more data types can implement. By declaring a type as an instance of a typeclass, one provides implementations of the typeclass’s methods for that particular type.
Typeclasses enable polymorphic functions to operate uniformly over all types that satisfy the typeclass interface. An example might be equality, such that one can use == on that data type.
Definition 3 (Type Saftey).
A guarantee that operations in a language will only be performed on compatible types, preventing type errors at compile time (or runtime).
Languages which are not type-safe may utilise typeannotation, where types for variables and functions are explicitly specifying to guide the compiler and improve code clarity, or type inference , where a compiler automatically deduces the types of expressions without explicit type annotations.
Type annotations are considered practically weaker because they serve mainly as optional documentation or hints rather than as fully enforced contracts. In many languages (e.g., Python), type annotations are not strictly checked at compile time or runtime. meaning that:
  • They can be incomplete or even outdated, lacking the automatic consistency guarantees that come with full type inference.
  • They rely on external tools (like linters or type checkers) rather than being integrated into the language’s core execution semantics.
  • The annotations do not prevent all type errors at runtime, as the language may still perform implicit conversions or delay error detection until execution.
Thus, while they improve code readability and can catch some errors, type annotations do not enforce type correctness as rigorously as a fully statically and automatically inferred type system.
@@ 1.5 Multigraphs
Multigraphs are not widely used as a standard computational representation in cheminformatics or chemical graph theory. Existing tools and frameworks generally rely on labeled graphs or weighted graphs to encode bond multiplicities and types.
@@
The set-theoretic definition of a multiset (or bag) is a generalization of a set where elements are allowed to appear more than once. Formally, a multiset $M$ over a universe $U$ is defined as a function:
@@ 1.6 Dietz representation
The constitution of a structure is represented in set theoretic notation, with two sets. The first set contains all atoms in the structure as triples, containing the number of unshared valence electrons of the atom, a numerical index, and the atomic symbol. The second set describes “bonding systems”, or bonding relationships between atoms, and contains pairs consisting of the number of shared valance electrons, and a set of edges representing bonds over which the electrons are shared.
@@
Delocalised bonding is represented by a bonding relationship containing multiple set of edges. For instance, the delocalised pi electrons of a benzene ring would be described by a bonding relationship of the form $(6,\{\{c_{1},c_{2}\},\{c_{2},c_{3}\},\{c_{3},c_{4}\},\{c_{4},c_{5}\},\{c_{5},c_{6}\},\{c_{6},c_{1}\}\})$, where $c_{i}$ is the numerical index of each carbon atom in the ring. The localised bonding between carbon atoms would be represented by $(2,\{\{c_{i},c_{j}\})$.
@@
There are two hydrogen atoms, labeled 1 and 2, each with zero unshared valance electrons in the atomic core, and a one covalent bond with two shared valance electrons.
@@
A major advantage of the Dietz representation is its general model for the representation of bonding. As a result, it is straightforward to describe complex bonding systems such as organometallics, with ferrocene and diborane being given as examples in the original article. The lack of labeling requirements for delocalised bonding stands in contrast to string based representations such as SMILES, which has explicit rules in the grammar (a form of labeling) for aromatic carbon, and two rules for different cases of aromatic nitrogen.
@@ 1.7 SMILES
Corresponding to molecular graphs, SMILES inherit the limitations of this representational framework. While SMILES can represent aromaticity through specific syntax[4], it cannot encode delocalized bonding beyond such cases. Resonance and tautomerism must either be simplified into a single structure or require multiple distinct structures to represent each form.
@@
Attempts to canonicalize SMILES, and the proprietary nature of the Daylight canonicalisation algorithm, have led to inconsistencies between implementations and research teams [12], though the OpenSMILES initiative has mitigated this issue by providing a standardized grammar.
@@ 1.8 SELFIES
Table 1: Derivation rules of SELFIES [5] for molecules in the QM9 dataset.
@@
Furthermore, E-Z stereochemistry does not in general provide accurate dihedral or torsional angles. Proteins rely on these angles for many tasks, which may be critical for machine learning tasks [29]. As Dayalan explains ([22]) “Dihedral angles are of considerable importance in protein structure prediction as they define the backbone of a protein, which together with side chains define the entire protein conformation.”. Angles influence the energy associated with a conformation of a drug which in turn relates to the utility and stability of the drug [30]. Although SELFIES solves the syntactic invalidity problem of SMILES, many challenges remain.
1.9 The future of SELFIES and of molecular representation - Extensions of SELFIES and future projects
@@ 1.11 Barriers to a general, digital, molecular representation
The lack of quantum chemical information, which is not currently part of any modern digital representation, is also a barrier to representation, as expresed in [5] article:
@@ 1.14 Extensions of SMILES
New syntaxes or grammars for string-based representation have been introduced in an attempt to represent different kinds of molecules not representable with only SMILES. For instance, there are many syntaxes from just one vendor called Daylight111www.daylight.com.
@@
Balsa, a fully-specified language designed as a subset of SMILES, has been introduced [34]. It aims to address the lack of a comprehensive SMILES specification and the resulting practical problems, such as the diversity of implementations, difficulties in standardization efforts, and challenges in developing extensions. Balsa provides a formal, machine-readable grammar that defines its syntax and describes semantics at three levels: constitution, configuration, and conformation. However, while Balsa improves specificity and compatibility, its compact representation may sacrifice some readability compared to more verbose alternatives. Additionally, like SMILES, Balsa strings can become complex and difficult to interpret for larger, more intricate molecules.
1.15 Molecular Fingerprints
Molecular fingerprints encode structural or chemical features of a molecule, typically as a vector, and are used for comparison, searching, and modeling in cheminformatics[35, 12].
Different structural features may be encoded depending on the fingerprinting method. Some fingerprints capture topological information, as seen in Atom Pair and Topological Torsion fingerprints, which encode molecular connectivity and atomic relationships based on graph traversal [35, 36]. Others, such as Extended Connectivity Fingerprints (ECFP) and Functional Class Fingerprints (FCFP), encode local atomic environments, iteratively expanding circular neighborhoods around each atom [36, 13]. Pharmacophore fingerprints, including Pharmacophore Pairs (PH2) and Pharmacophore Triplets (PH3), emphasize functional group interactions, encoding features related to molecular recognition rather than direct connectivity [36]. Some fingerprints, such as MinHashed Atom Pair (MAP4) and LINGO, operate directly on molecular strings, such as SMILES, instead of deriving features from a molecular graph [35, 36].
Although often described as vectors, the data type of fingerprint representations which use hashing functions do not formally meet the definition of a vector over a vector space.
Definition 8 (Vector Space).
A vector space over a field $\mathbb{F}$ is a set $V$ equipped with two operations:
  • Vector addition: A binary operation $+:V\times V\to V$ that assigns to every pair of vectors $u,v\in V$ a vector $u+v\in V$.
  • Scalar multiplication: A binary operation $\cdot:\mathbb{F}\times V\to V$ that assigns to every scalar $\lambda\in\mathbb{F}$ and every vector $v\in V$ a vector $\lambda v\in V$.
These operations must satisfy the following axioms for all $u,v,w\in V$ and all scalars $\lambda,\mu\in\mathbb{F}$:
  1. 1.
    (Associativity of addition) $(u+v)+w=u+(v+w)$.
  2. 2.
    (Commutativity of addition) $u+v=v+u$.
  3. 3.
    (Existence of an additive identity) There exists an element $0\in V$ such that $v+0=v$ for all $v\in V$.
  4. 4.
    (Existence of additive inverses) For every $v\in V$, there exists an element $-v\in V$ such that $v+(-v)=0$.
  5. 5.
    (Distributivity of scalar multiplication over vector addition) $\lambda(u+v)=\lambda u+\lambda v$.
  6. 6.
    (Distributivity of scalar multiplication over field addition) $(\lambda+\mu)v=\lambda v+\mu v$.
  7. 7.
    (Associativity of scalar multiplication) $(\lambda\mu)v=\lambda(\mu v)$.
  8. 8.
    (Existence of a multiplicative identity) There exists a scalar $1\in\mathbb{F}$ such that $1v=v$ for all $v\in V$.
In hashed fingerprints, such as MAP4, the hashing algorithm SH1 is used to distribute data uniformly over a range of outputs, in a way whic is non-invertible, non-continuous, and irreversible. Since SHA-1 and MinHashing introduce non-linearity in the MAP4 fingerprinting process, the transformation from molecular structure to MAP4 does not preserve vector space properties such as addition, scalar multiplication, or linear combination.
Consider the following hypothetical MAP4 fingerprints of molecules A and B:
$\mathbf{u}$ $=MAP4(A)=(101,203,405,512)$
$\mathbf{v}$ $=MAP4(B)=(98,207,410,515)$
Operations such as vector addition, or scalar multiplication are not well-defined, and it is unclear what additive or multiplicate identities, or additive inverses would represent. For instance, are the following meaningful operations in the vector space, and do they correspond to valid molecular fingerprints?
$\mathbf{u}+\mathbf{v}$ $=(101+98,203+207,405+410,512+515)$
$=(199,410,815,1027)\quad?$
[…]
$=(0,0,0,0)\quad?$
$2\mathbf{u}$ $=(2\times 101,2\times 203,2\times 405,2\times 512)$
$=(202,406,810,1024)\quad?$
The term “vector” may be a terminological convenience, but it is not an accurate reflection of the data type. It may well be detrimental to cross-field communication to see simple mathematical terminology, which has a rigorous definition, carelessly used. However, it appears that cheminformaticians are more concerned with computing the Tanimoto similarity [37] or the dot product between two vectors, rather than taking into account the full definition of a vector.